OpenStudy (anonymous):

Differentiate sin^2(cos * sq. root of (sin * pi * x) )

OpenStudy (agent0smith):

\[\LARGE \frac{ d }{ dx } \sin ^2 \left( \cos \sqrt{\sin(\pi x)} \right) = \]

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Ugh, I'm too hungry to do that right now... you have to use the chain rule a handful of times, but it's gonna get ugly since you've got functions within functions within functions.